# Enhancing ground-state interaction strength of neutral atoms via Floquet stroboscopic dynamics

### v1

@@ III Performance Assessment
III.1 Wide-Range Interaction Fidelity and Quasienergy Spectrum Analysis
@@
So far, the analysis has been based on idealized theoretical assumptions. To explore the feasibility of our scheme in broader scenarios and to examine its dependence on the Rydberg blockade effect, we analyze its performance under different Rydberg interaction strengths with two atoms. Figure 3(a) shows the fidelity distribution of the preparation of the state $|W_{2}\rangle$, with the vertical and horizontal axes representing the total number of pulse periods $N$ and the Rydberg interaction strength $U_{rr}$. Consequently, each vertical slice in the figure corresponds to a complete dynamical simulation with a different total evolution time $NT$. The figure reveals two key features, visible in both the full parameter scan (panel i) and the locally magnified region (panel ii). Firstly, in the vast majority of the parameter space, the scheme achieves fidelities that exceed $99\%$ (white regions). In the weak Rydberg interaction regime, the scheme still achieves efficient state preparation, demonstrating its robustness to reduced interaction strength. Secondly, a low-fidelity vertical canyon emerges in certain parameter regions, most prominently around $U_{rr}/\Omega\approx 1.25$, as shown in panel (ii). This systematic feature persists in both small and large $N$, indicating a strong dip in fidelity at this interaction strength. In addition, we conducted analyzes with three and four atoms in Appendix D, and the results were consistent with those obtained for two atoms.
@@
To understand the fidelity decrease in these regions, as shown in Fig. 3(b), we analyzed the quasienergy spectrum of the system. At the point where fidelity decreases most significantly, around $U_{rr}/\Omega\approx 1.25$, the quasienergies $E_{3}$, $E_{5}$, and $E_{6}$ are closely spaced, leading to pronounced nonadiabatic transitions during evolution, as shown in Fig. 3(b) and its zoomed-in view in Fig. 3(c). The corresponding eigenstates are $|\psi_{3}\rangle\approx 0.65|gg\rangle+0.72|W_{2}\rangle$, $|\psi_{5}\rangle\approx 0.71|gg\rangle+0.57|W_{2}\rangle$, and $|\psi_{6}\rangle\approx 0.16|gg\rangle+0.37|W_{2}\rangle+0.69|T_{2}\rangle+0.46|D_{2}\rangle+0.33|rr\rangle$. Ideally, the system should adiabatically follow the $|gg\rangle\rightarrow|W_{2}\rangle$ pathway. However, small energy gaps cause part of the population to be transferred from $|\psi_{3}\rangle$ and $|\psi_{5}\rangle$ to $|\psi_{6}\rangle$, which contains significant components $|ee\rangle$, $|D_{2}\rangle$, and $|rr\rangle$. This transition results in a noticeable enhancement of the $|ee\rangle$ population. For other regions where the fidelity does not exhibit a significant decrease, such as at $U_{rr}/\Omega\approx 2.58$, an avoided crossing occurs between $E_{5}$ and $E_{6}$, as shown in Fig. 3(b) and its zoomed-in view in Fig. 3(d). The corresponding eigenstates are $|\psi_{5}\rangle\approx 0.63|gg\rangle+0.68|W_{2}\rangle+0.29|rr\rangle$, $|\psi_{6}\rangle\approx 0.29|gg\rangle+0.18|W_{2}\rangle+0.18|ee\rangle+0.22|D_{2}\rangle+0.85|rr\rangle$. The reduced energy gap leads to strong mixing between the two eigenstates, significantly increasing the nonadiabatic transition probability in the avoided-crossing region. As a result, part of the population is transferred from $|\psi_{5}\rangle$ to $|\psi_{6}\rangle$, which has higher weights $|ee\rangle$ and $|rr\rangle$, thus further enhancing the population of $|ee\rangle$. The independence from the number of driving cycles $N$ confirms that this effect originates from an intrinsic energy matching condition of the Floquet Hamiltonian, rather than from a dynamical process dependent on timescale. Apart from the dark blue region, fidelity remains remarkably close to unity. Interestingly, even for values of the Rydberg interactions much smaller than those corresponding to the blockade regime, fidelity remains very large. Notice that for very small interaction strengths, we may maintain large fidelities by increasing the number of cycles. In contrast, most previous schemes strongly rely on the Rydberg blockade effect [37, 39, 36], whereas our method does not. Although the performance of the scheme is reduced when $U_{rr}/\Omega\approx 1.25$, its key advantage lies in the fact that, under periodic driving, fluctuations in atomic positions cause the Rydberg interaction strength to vary rather than remain fixed at a specific value during each dynamical evolution. The combined effect of periodic driving and positional fluctuations effectively mitigates the detrimental influence of unfavorable Rydberg interaction strengths. This feature makes our scheme particularly suitable for future implementations in atomic arrays with well-defined interatomic spacing [84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99].
Figure 4: (a) The fidelity of the state $|W_{\mathbb{N}}\rangle$ during evolution with and without decay from Rydberg state $|r\rangle$. We set the decay rate $\gamma/2\pi=1$ kHz for the room-temperature (300 K) case. Fidelity of the target state $|W_{\mathbb{N}}\rangle$ during the evolution for $\mathbb{N}=4$ atoms under different noise conditions: (b) Rabi amplitude noise, (c) laser phase error, and (d) laser detuning error. Other parameters are the same with Fig. 2.
@@ III.2 Robustness Against Experimental Noise
Building on the mechanism described above, we next extend the analysis from a two-atom system to a four-atom system and validate the scheme under realistic conditions by examining the impact of key experimental imperfections, including spontaneous emission from the Rydberg state $|r\rangle$, laser phase error, Rabi amplitude noise, and laser detuning error. We point out that although our numerical calculations are restricted to the case of $\mathbb{N}=4$ due to the large Hilbert space, the proposed scheme is equally applicable to larger systems ($\mathbb{N}>4$).
Firstly, to evaluate the performance of the protocol under more realistic conditions, we simulated the dynamics of systems by introducing a form of temporally quenched disorder. In a realistic physical system, the influence of the environment is unavoidable. Considering the spontaneous emission of Rydberg states, the effective Lindblad master equation is given by:
@@
$\dot{\rho}=\sum_{i=1}^{\mathbb{N}}\frac{\gamma}{2}\mathcal{D}[|e_{i}\rangle\langle r_{i}|]\rho+\frac{\gamma}{2}\mathcal{D}[|g_{i}\rangle\langle r_{i}|]\rho,$ (6)
where $\mathcal{D}[o]\rho=[o\rho o^{\dagger}-(o^{\dagger}o\rho+\rho o^{\dagger}o)/2]$ and $\gamma$ are the effective spontaneous emission rates. Based on the above scheme, we numerically simulated the fidelity $F(t)=\langle W_{\mathbb{N}}|\rho(t)|W_{\mathbb{N}}\rangle$ of the evolution using the master equation, as shown in Fig. 4(a). This is particularly relevant for atomic ensembles, where $U_{rr}$ can vary due to thermal motion or shot-to-shot positional variations. In our simulation, $U_{rr}$ was randomly resampled in each driving cycle from a uniform distribution between 0 and $2\pi\times 200$ MHz. Interestingly, despite this strong cycle-to-cycle disorder, the final fidelity of the state $|W_{\mathbb{N}}\rangle$ consistently exceeds $99.1\%$. Even considering the kHz-scale spontaneous decay rate of Rydberg states at room temperature (300 K), the fidelity remains remarkably high. The key insight is that, with cycle-to-cycle fluctuations, the system rarely falls exactly at $U_{rr}/\Omega\approx 1.25$. Thus, the impact of the resonant leakage channel is strongly suppressed. This demonstrates that the protocol is not only robust against static parameter offsets but also highly resilient to dynamic noise.
Secondly, due to the instability of the laser or the presence of environmental noise, random fluctuations in the driving field intensity result in Rabi amplitude noise of the laser [100, 101, 102]. The Rabi frequency of laser $\Omega$ in atom $i$ can be considered as $(1+\epsilon)\Omega$. Within a reasonable range of error, the maximum Rabi amplitude noise is set to $\epsilon=0.01$ as Gaussian random dynamical noise. We numerically simulated the fidelity $F(t)$ as shown in Fig. 4(b). The simulation results demonstrate that Rabi amplitude noise has a slight impact on the scheme, but the fidelity is still above $97.1\%$. In practical applications, due to the influence of quantum noise and other factors, the system deviates from an ideal state, and the laser phase often exhibits significant deviations that cannot be ignored. Usually, the laser phase error can be expressed as $\varphi(t)=\varphi_{0}+\delta\varphi(t)$. In the presence of errors, the Rabi frequency is adjusted to $\Omega e^{i\delta\varphi(t)}$. Within the error margin of the current experimental environment [103, 104, 105], we can assume that the laser phase noise is around $\delta\varphi=0.1\pi$. In the simulation process, we consider this error to be a random dynamic phase error of the laser, as shown in Fig. 4(c). The simulation results demonstrate that the phase error has a negligible impact on the scheme.
Finally, laser detuning is a critical error source in quantum optics and quantum computing, as it shifts the effective transition frequency and weakens the Rydberg blockade in Rydberg atoms. In Fig. 4(d), we simulate the evolution of fidelity under detuning errors ranging from 30-30 kHz to $30$ kHz. The selected detuning errors were randomly chosen within this range, and we found that the fidelity reached up to $98.8\%$ throughout this error spectrum. However, it is worth noting that the detuning error should not be excessively large to maintain such high fidelity. Additionally, this result highlights the robustness of the system within the specified range.

### v2

@@ III Performance Assessment
III.1 Fidelity and Quasienergy Spectrum Analysis
@@
So far, the analysis has been based on idealized theoretical assumptions. To explore the feasibility of our scheme in broader scenarios and to examine its dependence on the Rydberg blockade effect, we analyze its performance under different Rydberg interaction strengths with two atoms. Figure 3(a) shows the fidelity distribution of the preparation of the state $|W_{2}\rangle$, with the vertical and horizontal axes representing the total number of pulse periods $N$ and the Rydberg interaction strength $U_{rr}$. Consequently, each vertical slice in the figure corresponds to a complete dynamical simulation with a different total evolution time $NT$. The figure reveals two key features, visible in both the full parameter scan (panel i) and the locally magnified region (panel ii). Firstly, in the vast majority of the parameter space, the scheme achieves fidelities that exceed $99\%$ (white regions). More interestingly, in the weak Rydberg interaction regime, the scheme still achieves efficient state preparation, demonstrating its robustness to reduced interaction strength. Secondly, a low-fidelity vertical canyon emerges in certain parameter regions, most prominently around $U_{rr}/\Omega\approx 1.25$, as shown in panel (ii). This systematic feature persists in both small and large $N$, indicating a strong dip in fidelity at this interaction strength. In addition, we conducted analyzes with three and four atoms in Appendix D, and the results were consistent with those obtained for two atoms.
@@
To understand the fidelity decrease in these regions, as shown in Fig. 3(b), we analyzed the quasienergy spectrum of the system. At the point where fidelity decreases most significantly, around $U_{rr}/\Omega\approx 1.25$, the quasienergies $E_{3}$, $E_{5}$, and $E_{6}$ are closely spaced, leading to pronounced nonadiabatic transitions during evolution, as shown in Fig. 3(b) and its zoomed-in view in Fig. 3(c). The corresponding eigenstates are $|\psi_{3}\rangle\approx 0.65|gg\rangle+0.72|W_{2}\rangle$, $|\psi_{5}\rangle\approx 0.71|gg\rangle-0.57|W_{2}\rangle$, and $|\psi_{6}\rangle\approx 0.16e^{-0.14i}|gg\rangle+0.37e^{3.00i}|W_{2}\rangle+0.69|T_{2}\rangle+0.46e^{0.02i}|D_{2}\rangle+0.33e^{-2.41i}|rr\rangle$. Ideally, the system should adiabatically follow the $|gg\rangle\rightarrow|W_{2}\rangle$ pathway. However, small energy gaps induce transitions from $|\psi_{3}\rangle$ and $|\psi_{5}\rangle$ to $|\psi_{6}\rangle$, which contain significant components $|ee\rangle$, $|D_{2}\rangle$, and $|rr\rangle$. This transition results in a noticeable enhancement of the $|ee\rangle$ population. For other regions where the fidelity does not exhibit a significant decrease, such as at $U_{rr}/\Omega\approx 2.58$, an avoided crossing occurs between $E_{5}$ and $E_{6}$, as shown in Fig. 3(b) and its zoomed-in view in Fig. 3(d). The corresponding eigenstates are $|\psi_{5}\rangle\approx-0.63|gg\rangle+0.68|W_{2}\rangle+0.29e^{-1.62i}|rr\rangle$, $|\psi_{6}\rangle\approx 0.29e^{1.66i}|gg\rangle+0.18e^{-1.48i}|W_{2}\rangle+0.18e^{-1.36i}|ee\rangle+0.26e^{1.65i}|T_{2}\rangle+0.22e^{-1.50i}|D_{2}\rangle+0.85|rr\rangle$. The reduced energy gap leads to strong mixing between the two eigenstates, significantly increasing the nonadiabatic transition probability in the avoided-crossing region. As a result, part of the population is transferred from $|\psi_{5}\rangle$ to $|\psi_{6}\rangle$, which has higher weights $|ee\rangle$ and $|rr\rangle$, thus further enhancing the population of $|ee\rangle$. The independence from the number of driving cycles $N$ confirms that this effect originates from an intrinsic energy matching condition of the Floquet Hamiltonian, rather than from a dynamical process dependent on timescale. Apart from the dark blue region, fidelity remains remarkably close to unity. Interestingly, even for values of the Rydberg interactions much smaller than those corresponding to the blockade regime, the fidelity remains very large. Notice that for very small interaction strengths, we may maintain large fidelities by increasing the number of cycles. In contrast, most previous schemes strongly rely on the Rydberg blockade effect [38, 40, 37], whereas our method does not. Although the performance of the scheme is reduced when $U_{rr}/\Omega\approx 1.25$, its key advantage lies in the fact that, under periodic driving, fluctuations in atomic positions cause the Rydberg interaction strength to vary rather than remain fixed at a specific value during each dynamical evolution. The combined effect of periodic driving and positional fluctuations effectively mitigates the detrimental influence of unfavorable Rydberg interaction strengths. This feature makes our scheme particularly suitable for future implementations in atomic arrays with well-defined interatomic spacing [88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103].
Figure 4: Fidelity of the target state $|W_{\mathbb{N}}\rangle$ as a function of the cycle number under different noise and decoherence conditions. (a) Fidelity of the target state $|W_{\mathbb{N}}\rangle$ for $\mathbb{N}=4$ atoms in the presence of atomic position disorder, and the influence of additional decoherence effects including spontaneous decay, blackbody-radiation-induced decay, and intermediate-state scattering. (b–d) Fidelity for $\mathbb{N}=4$ atoms in the presence of different technical noises: (b) Rabi amplitude noise, (c) laser phase noise, and (d) laser detuning noise. (e) Fidelity for different atom numbers $\mathbb{N}=4,5,6$ in the presence of Doppler-induced dephasing at atomic temperatures of $5\,\mu\mathrm{K}$ and $12\,\mu\mathrm{K}$. (f) Same as panel (e), but with an alternating two-photon driving scheme in successive cycles to suppress the accumulated Doppler-induced dephasing. Other parameters are the same with Fig. 2.
@@ III.2 Robustness Against Experimental Noise
Building on the mechanism described above, we now extend the analysis from a two-atom system to a four-atom system and validate the scheme under realistic conditions by examining the impact of key experimental imperfections, including static positional disorder of atoms during the driving cycles, spontaneous emission from the Rydberg state $|r\rangle$, blackbody-radiation-induced decay, scattering via the intermediate state, laser phase error, Rabi amplitude noise, laser detuning error, as well as Doppler-induced frequency shifts. We point out that although our numerical calculations are restricted to the case of $\mathbb{N}=4$ due to the large Hilbert space, the proposed scheme is equally applicable to larger systems ($\mathbb{N}>4$).
Intrinsic atomic decay and disorder.— To evaluate the robustness of the protocol under realistic experimental conditions, we incorporate intrinsic decoherence and static disorder into the system dynamics. The dominant noise sources include spontaneous decay of the Rydberg state, blackbody-radiation-induced transitions among Rydberg levels, and effective dephasing arising from intermediate-state scattering during the two-photon excitation process. The resulting dynamics are described by the Lindblad master equation
@@
$\dot{\rho}=$ $\sum_{i=1}^{\mathbb{N}}\Bigg[\frac{\gamma}{2}\mathcal{D}\!\left[|e_{i}\rangle\langle r_{i}|\right]\rho+\frac{\gamma}{2}\mathcal{D}\!\left[|g_{i}\rangle\langle r_{i}|\right]\rho$
$\quad+\gamma_{\mathrm{BBR}}\mathcal{D}\!\left[|p_{i}\rangle\langle r_{i}|\right]\rho+\gamma_{\mathrm{IS}}\mathcal{D}\!\left[|r_{i}\rangle\langle r_{i}|\right]\rho\Bigg],$ (6)
where $\mathcal{D}[o]\rho=o\rho o^{\dagger}-(o^{\dagger}o\rho+\rho o^{\dagger}o)/2$ is the standard Lindblad superoperator. Here, the Rydberg state is $|r\rangle=|80S_{1/2}\rangle$, with spontaneous decay $\gamma/2\pi=0.28~\mathrm{kHz}$, blackbody-induced transitions $\gamma_{\mathrm{BBR}}/2\pi=0.50~\mathrm{kHz}$, and effective dephasing $\gamma_{\mathrm{IS}}$ from intermediate-state scattering. For the two-photon excitation scheme, the intermediate-state scattering rate is given by [73]
$\gamma_{\mathrm{IS}}=\frac{\Omega_{780}^{2}+\Omega_{480}^{2}}{4\Delta_{p}^{2}}\Gamma_{e},$ (7)
where $\Gamma_{e}=2\pi\times 6~\mathrm{MHz}$ is the natural linewidth of the intermediate state. The single-photon Rabi frequencies are $\Omega_{780}=2\pi\times 237~\mathrm{MHz}$ and $\Omega_{480}=2\pi\times 303~\mathrm{MHz}$, and the detuning is $\Delta_{p}=2\pi\times 7.8~\mathrm{GHz}$. Based on the above scheme, we first examine the impact of interaction inhomogeneity, which is particularly relevant for atomic ensembles where the van der Waals interaction strength $U_{rr}$ fluctuates due to thermal motion or shot-to-shot positional variations. In our simulations, $U_{rr}$ is randomly sampled in each driving cycle from a uniform distribution over $0,\,2\pi\times 200~\mathrm{MHz}]$. The resulting fidelity, $F(t)=\langle W_{\mathbb{N}}|\rho(t)|W_{\mathbb{N}}\rangle$, is shown in Fig. 4(a). Remarkably, despite this strong cycle-to-cycle disorder, the final fidelity of the target state $|W_{\mathbb{N}}\rangle$ consistently exceeds $99.1\%$. We next perform a more comprehensive simulation that simultaneously incorporates all dominant decoherence mechanisms at room temperature ($300~\mathrm{K}$), including spontaneous decay of the Rydberg state, blackbody-radiation-induced transitions, and scattering via the intermediate state involved in the off-resonant two-photon excitation. The intermediate-state scattering is modeled using the standard treatment of off-resonant Raman processes [9, 104, 73]. Even under the combined influence of these dissipation channels, the fidelity of the target state remains as high as $97.5\%$. This robustness can be attributed to the cycle-to-cycle fluctuations of $U_{rr}$, which prevent the system from residing near the resonant condition $U_{rr}/\Omega\approx 1.25$ for extended durations, thereby strongly suppressing leakage through resonant channels. These results demonstrate that the protocol is resilient against both intrinsic decoherence and realistic dynamical disorder encountered in experiments.
Laser Rabi amplitude noise and phase error.— In realistic experiments, imperfections in the driving fields play an important role. Laser instability and technical noise induce stochastic fluctuations in the driving-field intensity, which can be effectively modeled as Rabi-amplitude noise [105, 106, 107].Accordingly, the Rabi frequency acting on atom $i$ is taken as $\Omega_{i}=[1+\epsilon]\Omega$. Within a realistic error budget, the Rabi-amplitude noise $\epsilon$ is modeled as a random variable following the Gaussian distribution $\mathcal{N}(0,\sigma_{\epsilon}^{2})$ with a standard deviation of $\sigma_{\epsilon}=0.01$ (corresponding to $1\%$ amplitude noise), representing a noise level that is well within the capabilities of current experimental amplitude-control techniques. By averaging over 50 independent noise realizations, we obtain the fidelity $F(t)$ shown in Fig. 4(b). The results show that the protocol is only weakly affected by Rabi-amplitude noise, with the final fidelity remaining above $97.1\%$. We further consider laser phase noise, which arises from both quantum noise and technical fluctuations and leads to deviations from an ideal coherent drive. The laser phase can be written as $\varphi(t)=\varphi_{0}+\delta\varphi(t)$, resulting in a modified Rabi frequency $\Omega\rightarrow\Omega e^{i\delta\varphi(t)}$. Within the error margins reported in current experiments [108, 109, 110], a typical phase fluctuation of $\delta\varphi=0.1\pi$ is assumed. In the simulations, this effect is modeled as a random dynamical phase error, and the results are obtained by averaging over 50 independent noise realizations. The corresponding fidelity is shown in Fig. 4(c). It is found that such phase noise has a negligible impact on the protocol performance.
[…]
Doppler dephasing.— Doppler shifts associated with the coupling between the ground states $|g\rangle$ and $|e\rangle$ are negligible, as the coupling is mediated either by microwave fields with long wavelengths or by co-propagating optical fields with nearly identical wave vectors, resulting in a strongly suppressed effective wave vector. In contrast, for the two-photon excitation to the Rydberg state, the Doppler shift acts as an effective random detuning in a periodically driven scheme, which accumulates from cycle to cycle and leads to phase errors in the collective dynamics. This effect is incorporated by introducing an additional error Hamiltonian $H_{D}=\sum_{i=1}^{\mathbb{N}}\delta_{i}^{D}\,|r_{i}\rangle\langle r_{i}|$, where $\delta_{i}^{D}$ denotes the Doppler-induced detuning experienced by the $i$th atom during the two-photon Rydberg excitation. To account for atomic motion, each atom is assigned a velocity randomly sampled from a one-dimensional Maxwell–Boltzmann distribution corresponding to a finite atomic temperature $T_{a}$. The Doppler shift is given by $\delta_{i}^{D}=k_{\mathrm{eff}}v_{\mathrm{rms}}$, where $k_{\mathrm{eff}}=|k_{780}-k_{480}|$ denotes the effective wave vector along the excitation ($z$) axis, determined by the difference in magnitude between the wave vectors of the counter-propagating fields.. The root-mean-square atomic velocity along this direction is $v_{\mathrm{rms}}=\sqrt{k_{B}T_{a}/m}$, where $T_{a}$ is the atomic temperature and $m$ is the atomic mass. We consider an atomic ensemble confined in a geometry where the motion transverse to the excitation axis is tightly constrained and can be neglected. Under this approximation, the dominant motional fluctuations arise from atomic velocities projected along the excitation ($z$) direction, which can be described as thermal velocity fluctuations superimposed on a slowly varying center-of-mass motion. Such one-dimensional treatments of atomic motion are commonly used in studies of Doppler dephasing and collective effects in cold atomic ensembles [111]. The Doppler-induced collective dephasing of the $|W_{\mathbb{N}}\rangle$ state is shown in Fig. 4(e) for $\mathbb{N}=4,5,6$ atoms at different temperatures. Each data point is obtained by averaging over 200 independent numerical realizations of atomic velocities drawn from the corresponding thermal distribution. For $T_{a}=5~\mu\mathrm{K}$, the final fidelity is reduced to approximately $96.1\%$, while increasing the temperature to $T_{a}=12~\mu\mathrm{K}$ further lowers the fidelity to approximately $91.6\%$, reflecting the cumulative nature of Doppler dephasing under periodic driving. In view of the Doppler-induced degradation discussed above, we employ an alternating driving protocol, shown in Fig. 4(f), to reduce Doppler-induced effects. Reversing the propagation directions of the excitation beams in successive cycles flips the sign of the effective wave vector $k_{\mathrm{eff}}$, which causes the Doppler phase accumulated in one cycle to be partially compensated in the next. Consequently, long-term Doppler-induced dephasing is strongly suppressed, and the fidelity remains close to $99.0\%$ even at atomic temperatures up to $T_{a}=12~\mu\mathrm{K}$, demonstrating the strong robustness of the proposed protocol against Doppler noise.
III.3 Comparison of our scheme to previous schemes
Table 1: Comparison of our scheme to previous schemes
Reference Mechanism System Fidelity
Our scheme Floquet dynamics atomic ensemble $99\%$ (Theo.)
Ref. [55] Rydberg dressing atomic array $81\%$ (Exp.)
Ref. [56] Rydberg-antiblockade atomic array, $99\%$ (Theo.)
Ref. [57] induced dipole blockade regular chain $96\%$ (Theo.)
In this section, we situate our work within the broader context of neutral-atom quantum logic by comparing our Floquet stroboscopic dynamics with several representative schemes that realize effective ground-state blockade through Rydberg-mediated mechanisms, including Rydberg dressing [55], Rydberg antiblockade [56], and light-induced dipole blockade [57] as shown in Table 1.
First, the Rydberg dressing has been experimentally demonstrated as a powerful method for engineering effective interactions within the ground-state manifold [55]. In this approach, ground states are weakly and off-resonantly coupled to Rydberg states, producing tunable soft-core interactions. Nevertheless, a fundamental limitation of dressing schemes is the unavoidable admixture of the Rydberg character into the dressed states. As a result, the system is subject to continuous decoherence from the finite Rydberg lifetime, leading to an intrinsic trade-off between interaction strength and coherence time.
Subsequently, Ref. [56] proposed a scheme based on the Rydberg antiblockade mechanism [56], in which the interaction-induced energy shift is tuned to compensate for the laser detuning. This effectively maps the blockade physics entirely onto the ground-state manifold, without requiring the population of real excited states, and endows the protocol with a certain degree of robustness. However, the principal limitation of this approach lies in its poor scalability rather than its sensitivity to noise. Because the van der Waals interaction scales as $1/R^{6}$, the antiblockade condition can only be satisfied in highly constrained geometries with well-controlled interatomic separations. Consequently, the scheme is restricted to small planar atomic arrays and has been demonstrated only for systems of up to three atoms, making it difficult to extend to larger atomic systems or ensembles.
Finally, Ref. [57] investigated collective blockade effects mediated by dipole–dipole interactions. While this mechanism can induce photon or excitation blockade in dense atomic ensembles, it requires all atoms to be confined within a subwavelength volume ($R<\lambda/2\pi$). Such a stringent geometric constraint severely limits scalability and is incompatible with standard optical tweezer architectures, where individual addressability and readout typically require interatomic spacings on the order of micrometers.
In contrast, our Floquet-based protocol provides a robust and versatile route to realizing ground-state blockade with several key advantages. First, this ground-state blockade requires neither a predefined blockade radius nor strong vdW interactions, and it functions effectively without an additional cavity. Even in the regime of relatively weak Rydberg-Rydberg interactions, the effective ground-state blockade can still be established through Floquet engineering, substantially relaxing constraints on interatomic spacing and interaction strength. Second, the protocol is intrinsically robust against a wide range of experimental imperfections, including spontaneous emission, blackbody-radiation-induced decay, intermediate-state scattering, laser amplitude and phase noise, detuning errors, and Doppler-induced dephasing, as demonstrated by our numerical simulations. Finally, the blockade is effectively realized within the ground-state manifold at stroboscopic times, such that the system dynamics are well projected back onto the ground-state subspace at the end of each driving cycle. This feature effectively suppresses decoherence channels associated with finite Rydberg lifetimes and enables high-fidelity state preparation compatible with scalable, micrometer-spaced neutral-atom architectures.